OpenStudy (anonymous):

How to Define the following functions? help please.. 3. f(x)=x^2+1; g(x)=3x-2 a. f+g b. f-g c. f⋅g d. f/g e. g/f

OpenStudy (anonymous):

Part A: Add x^2 + 1 and 3x - 2 together and simplify. Part B: Subtract 3x - 2 from x^2 + 1 and simplify Part C: Multiply .......... You get the procedure now probably

OpenStudy (anonymous):

You won't be able to simplify d and e so don't stress too much on those.

OpenStudy (anonymous):

already have answer in A & B but for C,D,& E still dont have

OpenStudy (anonymous):

Part C : fg(x) = (x^ + 1)(3x - 2) = now just FOIL that out.

OpenStudy (anonymous):

please help how to solve C,D, & E

OpenStudy (anonymous):

C. f.g= (x^2+1).(3x-2) i use FOIL to get this (x*3x)+(x*-2)+(1*3x)+(1*-2) 3x^2+(-2x)+3x+(-2) f.g= 3x^2+x-2

OpenStudy (anonymous):

please check my answer...

OpenStudy (anonymous):

@mebs pls help here..

OpenStudy (anonymous):